OpenStudy (anonymous):

the sales tax on a $100 purchase is $15. how much would the sales tax be on a $40 purchase? Must use a proportion to solve.

OpenStudy (mathstudent55):

sales tax/purchase = 15/100 = x/40 Solve the proportion 15/100 = x/40 for x. Cross multiply to get started.

OpenStudy (anonymous):

What numbers do I multiply?

OpenStudy (mathstudent55):

Math is meant to be done on paper using proper fractions with proper horizontal fraction lines, not slashes. See pic below.